# Topological piezomagnetic effect in two-dimensional Dirac quadrupole altermagnets

### v1

@@ Appendix A Details of the orbital altermagnet model
where we can write $\sin k_{D}=2\sin(k_{D}/2)\cos(k_{D}/2)$. Now using that $\sin(k_{D}/2)=\sqrt{\delta/2t_{1}}$ one finds $\sin k_{D}=2\sqrt{(\delta/2t_{1})(1-\delta/2t_{1})}$. Defining the velocities $v_{1}$ and $v_{2}$ as
@@
$v_{2}$ $\equiv 4t_{d}\sin k_{D}=\frac{4t_{d}}{t_{1}}\sqrt{\delta(2t_{1}-\delta)},$ (13)
we obtain the Dirac Hamiltonian given by Eq. (2) for the valleys $\bf K$$_{\pm}$. The Dirac Hamiltonian for the valleys $\bf K$$^{\prime}_{\pm}$ is obtained directly by exchanging $q_{x}$ and $q_{y}$, which gives
$\mathcal{H}_{{\bf K}^{\prime}_{\pm}}=\pm(v_{1}q_{x}\tau^{z}+v_{2}q_{y}\tau^{x})+\Delta\tau^{y}.$ (14)
We thus obtain four gapped Dirac Hamiltonians for each of the four valleys, with $\Delta$ playing the role of the Dirac mass (and thus size of the gap).
Using this continuum model expansion, it is straightforward to compute the Berry curvature $\Omega_{xy}$ associated with the occupied valence band for each valley. We find the Berry curvatures of the different valleys as
@@
To compute the piezomagnetic polarizability within the continuum model, we also require an expansion of the strain term $W_{\bf k}$, for which, as in the main text, we only consider the coupling $\chi_{\bf k}=-2t_{0}(\cos k_{x}-\cos k_{y})$. The lowest order contribution to an expansion of the function $\chi_{\bf k}$ is a constant term (i.e., the value of $\chi_{\bf k}$ at $\bf K$$_{\pm}$ and $\bf K$$^{\prime}_{\pm}$), and we find
$\chi_{{\bf K}_{\pm}}$ $=-t_{0}(1-\cos k_{D})=-2t_{0}\sin^{2}\frac{k_{D}}{2}=-\frac{t_{0}\delta}{t_{1}},$ (18)
$\chi_{{\bf K}^{\prime}_{\pm}}$ $=-t_{0}(\cos k_{D}-1)=\frac{t_{0}\delta}{t_{1}}.$ (19)
Therefore, denoting $\mathcal{W}_{{\bf K}_{\pm}}$ and $\mathcal{W}_{{\bf K}^{\prime}_{\pm}}$ as the expansion of the strain perturbation (similar to $\mathcal{H}_{{\bf K}_{\pm}}$ and $\mathcal{H}_{{\bf K}^{\prime}_{\pm}}$), we have
$\mathcal{W}_{{\bf K}_{\pm}}=-w^{D}_{0},\qquad\mathcal{W}_{{\bf K}^{\prime}_{\pm}}=w^{D}_{0},$ (20)
with $w^{D}_{0}$ given by $w^{D}_{0}\equiv t_{0}\delta/t_{1}$.
Generalization of strain coupling. In the main text we have chosen the strain coupling simply as $W_{\bf k}=\chi_{\bf k}=-2t_{0}(\cos k_{x}-\cos k_{y})$, which amounts t introducing an anisotropy in the $t_{0}$ hopping with the appropriate $d_{x^{2}-y^{2}}$ symmetry. As defined in Eq. (3), the most general form of the strain coupling is $W_{\bf k}=\chi_{\bf k}+{\bf w}_{\bf k}\cdot\bm{\tau}$ and here we briefly comment on additional symmetry-allowed strain couplings that enter via ${\bf w}_{\bf k}$. The requirement is that all terms in $W_{\bf k}$ have the symmetry of the applied strain, which is of $d_{x^{2}-y^{2}}$ type. The chosen form of $\chi_{\bf k}$ clearly satisfies this requirement. An additional term which satisfies this requirement is given by
@@
Lattice model calculation of polarizability. In the lattice model calculation of the piezomagnetic polarizability discussed in the main text we evaluate Eq. (4) numerically. The integrand may be obtained analytically, however, and we find
$\Lambda=-\frac{2t_{0}e}{\hbar}\int\frac{d^{2}{\bf k}}{(2\pi)^{2}}\frac{t_{1}t_{d}\Delta(c_{x}-c_{y})^{2}(1-c_{x}c_{y})}{(\Delta^{2}+t^{2}_{x,{\bf k}}+t^{2}_{z,{\bf k}})^{3/2}}.$ (22)
Here we have abbreviated $c_{j}\equiv\cos k_{j}$.
When the more general form of the strain coupling is chosen [i.e., including $w_{x,{\bf k}}$ of Eq. (21)] the full expression for the polarizability $\Lambda$ is Venderbos (2025)
We have verified that computing $\Lambda$ with the more general form of the strain coupling does not affect the discontinuity at $\Delta=0$, and therefore does not affect the topological response.
Continuum model calculation of polarizability. A calculation of the polarizability within the continuum model is achieved in a straightforward way by applying Eq. (4) to each of the Dirac valleys and then summing the result. With the Berry curvature of the Dirac points given by Eqs. (15) and (16), and with the strain coupling given by Eq. (20), one finds
@@
$=-\frac{e}{\pi\hbar}w^{D}_{0}\text{sgn}(\Delta).$ (24)
Appendix B Continuum Dirac quadrupole model of the Lieb lattice altermagnet
@@
We will proceed as follows. First, we determine the continuum Dirac model by expanding the lattice Hamiltonian in the absence of strain, and then perform a similar expansion of the strain terms, for which we only keep the lowest zeroth order contributions. Based on the linearized continuum model of the full Hamiltonian, we then show in Appendix C how the piezomagnetic polarizability is computed.
@@
These function capture anisotropic next-nearest neighbor (and intra-sublattice) hoppings allowed in the presence of strain. In the continuum model expansion we only require the lowest order terms, which are constant $\bf q$}-independent terms obtained from evaluating $\chi_{\bf k}$ and $w_{z,{\bf k}}$ at $\bf K$$_{\pm}$ and $\bf K$$^{\prime}_{\pm}$. We find
@@
Appendix C Calculation of piezomagnetic polarizability for the Lieb lattice altermagnet
@@
For the valley $\bf K$$_{+}$, we can replace $\chi_{\bf k}$ by $w_{0}$ (see Appendix B) and we can take the vector ${\bf n}={\bf n}_{\bf q}$ as ${\bf n}=(-v_{1}q_{x}.-\zeta m,v_{2}q_{y})$. The vector $\bf w$} is simply given by ${\bf w}=(0,0,w^{D}_{z})$. With these ingredients we evaluate both the geometric and the interband terms in (LABEL:app:dM_a-Lieb) (i.e., the first and second terms in the integrand). From simple inspection we see that the geometric term is proportional to the Berry curvature of the occupied valence band, which is multiplied by the energy shift $w^{D}_{0}$. In the special case of the linearized Dirac model, the interband term can also be rewritten in terms of the Berry curvature (by multiplying and dividing by $v_{2}$), while $\partial_{y}\varepsilon_{{\bf K}_{+}+{\bf q}}$ becomes $\partial_{y}(-v_{0}q_{y})=-v_{0}$. Putting everything together, one obtains
@@
Appendix D Topological transition in the Lieb lattice altermagnet

### v2

@@ Appendix A Linearized continuum model for the orbital altermagnet
where we can write $\sin k_{D}=2\sin(k_{D}/2)\cos(k_{D}/2)$. Now using that $\sin(k_{D}/2)=\sqrt{\delta/2t_{1}}$ one finds $\sin k_{D}=2\sqrt{(\delta/2t_{1})(1-\delta/2t_{1})}$. It is then natural to define the two Dirac velocities $v_{1}$ and $v_{2}$ as
@@
$v_{2}$ $\equiv 4t_{d}\sin k_{D}=\frac{4t_{d}}{t_{1}}\sqrt{\delta(2t_{1}-\delta)}.$ (12)
Next, we expand the function ${\varepsilon}_{{\bf k}}$, which multiplies the identity. We find
${\varepsilon}_{{\bf K}_{\pm}+{\bf q}}$ $\approx-2t_{0}\cos^{2}\frac{k_{D}}{2}\pm t_{0}\sin k_{D}q_{y},$ (13)
${\varepsilon}_{{\bf K}^{\prime}_{\pm}+{\bf q}}$ $\approx-2t_{0}\cos^{2}\frac{k_{D}}{2}\pm t_{0}\sin k_{D}q_{x},$ (14)
and this motivates defining the velocity $v_{0}$ as
$v_{0}\equiv t_{0}\sin k_{D}=\frac{t_{0}}{t_{1}}\sqrt{\delta(2t_{1}-\delta)}$ (15)
and the Dirac point energy ${\varepsilon}_{D}$ as ${\varepsilon}_{D}\equiv-2t_{0}\cos^{2}(k_{D}/2)=-t_{0}(2-\delta/t_{1})$. The Dirac Hamiltonian for the valleys $\bf K$$_{\pm}$ is then given by (ignoring the constant term ${\varepsilon}_{D}$)
$\mathcal{H}_{{\bf K}_{\pm}}=\pm v_{0}q_{y}\pm(v_{1}q_{y}\tau^{z}+v_{2}q_{x}\tau^{x})+\Delta\tau^{y}.$ (16)
The Dirac Hamiltonian for the valleys $\bf K$$^{\prime}_{\pm}$ is obtained directly by exchanging $q_{x}$ and $q_{y}$. These Dirac Hamiltonians describe a gapped, anisotropic, and tilted (velocity $v_{0}$) Dirac cone. The gap is due to the mass term $\Delta$ and the anisotropy is a result of the generically different velocities $v_{1}$ and $v_{2}$. The tilt is caused by $\pm v_{0}q_{y}$. In the main text we have ignored this tilt due to $v_{0}$, since it does not enter the expression for the piezomagnetic polarizability when the strain coupling is chosen as $W_{\bf k}=-t_{0}(\cos k_{x}-\cos k_{y})$.
The continuum model thus consists of four gapped Dirac Hamiltonians for each of the four valleys. We note that, in the notation introduced in Eq. (3) of the main text, the Dirac Hamiltonian for the valleys $\bf K$$_{\pm}$ can be expressed as
$\mathcal{H}_{{\bf K}_{\pm}}={\varepsilon}^{{\bf K}_{\pm}}_{\bf q}+{\bf n}^{{\bf K}_{\pm}}_{\bf q}\cdot\bm{\tau},$ (17)
with
${\varepsilon}^{{\bf K}_{\pm}}_{\bf q}=\pm v_{0}q_{y},\quad{\bf n}^{{\bf K}_{\pm}}_{\bf q}=(\pm v_{2}q_{x},\Delta,\pm v_{1}q_{y}).$ (18)
A similar expression holds for the valleys $\bf K$$^{\prime}_{\pm}$.
Berry curvature. Using this continuum model expansion, it is straightforward to compute the Berry curvature $\Omega_{xy}$ associated with the occupied valence band for each valley. We find the Berry curvatures of the different valleys as
@@
Appendix B Generalized strain coupling in the orbital altermagnet
In the main text we have considered a particular choice of strain coupling, namely we have chosen $W_{\bf k}$ proportional to the identity: $W_{\bf k}=\chi_{\bf k}=-t_{0}(\cos k_{x}-\cos k_{y})$. In this Appendix we present the general form of $W_{\bf k}$ allowed by symmetry, i.e., a form for which ${\bf w}_{\bf k}$ is nonzero, and discuss how it affects the calculation of the orbital piezomagnetic polarizability.
Generalization of strain coupling. The orbital altermagnet model describes a band inversion of an $s$ and a $d$ band. In general, strain affects each band separately and leads to independent anisotropic dispersions of the $s$ and $d$ bands. This is captured by a strain perturbation $W_{\bf k}$ given by
$\chi_{\bf k}=-t_{a}(c_{x}-c_{y}),\quad w_{z,{\bf k}}=-t_{b}(c_{x}-c_{y})$ (22)
Here we have abbreviated $c_{j}\equiv\cos k_{j}$ and $t_{a}$ and $t_{b}$ are introduced as independent hopping amplitudes (different from $t_{0}$ and $t_{1}$) describing the effect of strain. To see how this relates to anisotropic hopping within the $s$ and a $d$ bands, we first note that nearest neighbor hopping amplitudes between the $s$ and $d$ states are given by $t_{s}=(t_{0}+t_{1})/2$ and $t_{d}=(t_{0}-t_{1})/2$, respectively. In the presence of strain, described by the (dimensionless) strain field $\phi$, these hoppings change as
$t_{s}=\frac{1}{2}(t_{0}+t_{1})$ $\rightarrow t_{s}(1\pm\eta_{s}\phi)$ (23)
$t_{d}=\frac{1}{2}(t_{0}-t_{1})$ $\rightarrow t_{d}(1\pm\eta_{d}\phi)$ (24)
where $\pm$ corresponds to the hopping along the $x/y$ direction, in accordance with an applied strain of ${\varepsilon}_{xx}-{\varepsilon}_{yy}$ symmetry. Here $\eta_{s}$ and $\eta_{d}$ are dimensionless parameters describing the generically different response of the $s$ and $d$ band hoppings to applied strain. It is then useful to define
$\eta_{\pm}\equiv\frac{1}{2}(\eta_{s}\pm\eta_{d}),$ (25)
such that the hoppings $t_{a}$ and $t_{b}$ introduced in Eq. (22) can be expressed in terms of $t_{0}$ and $t_{1}$ as
$t_{a}$ $=\eta_{+}t_{0}+\eta_{-}t_{1},$ (26)
$t_{b}$ $=\eta_{+}t_{1}+\eta_{-}t_{0}.$ (27)
Alternatively, one may express $\eta_{\pm}$ in terms of the hopping amplitudes $t_{0,1}$ and $t_{a,b}$ as
$\eta_{+}=\frac{t_{0}t_{a}-t_{1}t_{b}}{t^{2}_{0}-t^{2}_{1}},\quad\eta_{-}=\frac{t_{0}t_{b}-t_{1}t_{a}}{t^{2}_{0}-t^{2}_{1}}.$ (28)
Based on this discussion, the strain perturbation of Eq. (22) can be expressed in terms of $t_{0,1}$ and $\eta_{\pm}$ as
$\chi_{\bf k}$ $=-(\eta_{+}t_{0}+\eta_{-}t_{1})(\cos k_{x}-\cos k_{y}),$ (29)
$w_{z,{\bf k}}$ $=-(\eta_{+}t_{1}+\eta_{-}t_{0})(\cos k_{x}-\cos k_{y}),$ (30)
which represents the most general way in which strain can modify the hopping amplitudes of the $s$ and $d$ bands.
Further generalization. As defined in Eq. (3), the most general form of the strain coupling in the two-band orbital altermagnet model is $W_{\bf k}=\chi_{\bf k}+{\bf w}_{\bf k}\cdot\bm{\tau}$. Above we determined the form of $\chi_{\bf k}$ and $w_{z,{\bf k}}$, which correspond to intra-band strain terms. Here we briefly comment on additional symmetry-allowed strain interband couplings that enter via ${\bf w}_{\bf k}$, in particular $w_{x,{\bf k}}$. The requirement is that all terms in $W_{\bf k}$ have the symmetry of the applied strain, which is of $d_{x^{2}-y^{2}}$ type. An additional term which satisfies this requirement is given by
@@ Appendix B Generalized strain coupling in the orbital altermagnet
Appendix C Piezomagnetic polarizability of the orbital altermagnet
In this Appendix we describe in more detail how to calculate the orbital piezomagnetic polarozability given by Eq. (4) of the main text. We first briefly comment on the full lattice model calculation. Then we determine how the strain coupling enters in the continuum model formulation. This serves as a basis for calculating the piezomagnetic polarizability within the continuum Dirac model.
Lattice model calculation of polarizability. In the lattice model calculation of the piezomagnetic polarizability discussed in the main text, we evaluate Eq. (4) numerically using the full lattice tight-binding model. Choosing the strain coupling as $W_{\bf k}=\chi_{\bf k}+w_{z,{\bf k}}\tau^{z}$, with $\chi_{\bf k}$ and $w_{z,{\bf k}}$ given by Eqs. (29) and (30), the integrand of Eq. (4) may be obtained analytically, and we find
$\Lambda=\frac{2e\eta_{-}(t^{2}_{0}-t^{2}_{1})}{\hbar}\int\frac{d^{2}{\bf k}}{(2\pi)^{2}}\frac{t_{d}\Delta(c_{x}-c_{y})^{2}(1+c_{x}c_{y})}{(\Delta^{2}+t^{2}_{x,{\bf k}}+t^{2}_{z,{\bf k}})^{3/2}}.$ (32)
Here we have abbreviated $c_{j}\equiv\cos k_{j}$. Note that, based on Eq. (28), we may also write $\eta_{-}(t^{2}_{0}-t^{2}_{1})=t_{0}t_{b}-t_{1}t_{a}$. The results shown in Fig. 2(b) are obtained from numerically evaluating Eq. (32) for the case $t_{b}=0$ and $t_{a}=t_{0}$.
Strain coupling in the continuum model. To compute the piezomagnetic polarizability within the continuum model, we require an expansion of the strain term $W_{\bf k}$. To obtain such an expansion we proceed as in Appendix A. The lowest order contribution to an expansion of the function $\chi_{\bf k}$ is a constant term (i.e., the value of $\chi_{\bf k}$ at $\bf K$$_{\pm}$ and $\bf K$$^{\prime}_{\pm}$), and we find
$\chi_{{\bf K}_{\pm}}$ $=-(\eta_{+}t_{0}+\eta_{-}t_{1})(1-\cos k_{D}),$ (33)
$=-(\eta_{+}t_{0}+\eta_{-}t_{1})\delta/t_{1}=-\chi_{{\bf K}^{\prime}_{\pm}}.$ (34)
Similarly, the lowest order contribution to $w_{z,{\bf k}}$ is found as
$w_{z,{\bf K}_{\pm}}$ $=-(\eta_{+}t_{1}+\eta_{-}t_{0})(1-\cos k_{D}),$ (35)
$=-(\eta_{+}t_{1}+\eta_{-}t_{0})\delta/t_{1}=-w_{z,{\bf K}^{\prime}_{\pm}}.$ (36)
As a result, denoting the expansion of the strain perturbation as $\mathcal{W}_{{\bf K}_{\pm}}$ and $\mathcal{W}_{{\bf K}^{\prime}_{\pm}}$ (in analogy with $\mathcal{H}_{{\bf K}_{\pm}}$ and $\mathcal{H}_{{\bf K}^{\prime}_{\pm}}$), we have
$\mathcal{W}_{{\bf K}_{\pm}}=-w^{D}_{0}-w^{D}_{z}\tau^{z},\qquad\mathcal{W}_{{\bf K}^{\prime}_{\pm}}=w^{D}_{0}+w^{D}_{z}\tau^{z},$ (37)
with $w^{D}_{0}$ and $w^{D}_{z}$ defined as
$w^{D}_{0}$ $\equiv(\eta_{+}t_{0}+\eta_{-}t_{1})\delta/t_{1}=t_{a}\delta/t_{1},$ (38)
$w^{D}_{z}$ $\equiv(\eta_{+}t_{1}+\eta_{-}t_{0})\delta/t_{1}=t_{b}\delta/t_{1}.$ (39)
Here we have used Eqs. (26) and (27). The expansion of (37) is the basis for determining $\Lambda$ within the continuum model. We note in passing that (31), if included in the strain coupling, does not enter the continuum model and thus does not contribute to the polarizability (at this level).
Continuum model calculation of polarizability. A calculation of the polarizability within the continuum model is achieved in a straightforward way by applying Eq. (4) to each of the Dirac valleys and then summing the result. Here we proceed in two steps: we first determine the orbital piezomagnetic polarizability for the strain coupling considered in the main text, which makes the choice $t_{a}=t_{0}$ and $t_{b}=0$, and then determine the polarizability for the general case. As noted, in this case only the first term (i.e., “geometric” term) in Eq. (4) contributes and the polarizability directly follows from the Berry curvature and energy shifts. In this restricted case, with the Berry curvature of the Dirac points given by Eqs. (19) and (20), and with the strain coupling given by Eq. (38) (and $t_{a}$ set to $t_{0}$), one finds
@@ Appendix C Piezomagnetic polarizability of the orbital altermagnet
$=-\frac{e}{\pi\hbar}w^{D}_{0}\text{sgn}(\Delta)=-\frac{e}{\pi\hbar}(t_{0}\delta/t_{1})\text{sgn}(\Delta).$ (40)
This is the result quoted and discussed in the main text.
In the case of the more general strain coupling given by Eqs. (29) and (30), and in the continuum model by Eqs. (29) and (30), the second term in Eq. (4) also contributes to $\Lambda$. A key observation is that for the continuum Dirac model the contribution from the second term can still be expressed in terms of the Berry curvature of the Dirac points. Specifically, for the valleys $\bf K$$_{\pm}$ we find that the second term gives a contribution given by
$(\partial_{y}{\varepsilon}^{{\bf K}_{\pm}}_{\bf q})\frac{{\bf w}\cdot{\bf n}^{{\bf K}_{\pm}}_{\bf q}\times\partial_{x}{\bf n}^{{\bf K}_{\pm}}_{\bf q}}{2\left|{\bf n}^{{\bf K}_{\pm}}_{\bf q}\right|^{3}}\rightarrow w^{D}_{z}\frac{v_{0}}{v_{1}}\Omega^{{\bf K}_{\pm}}_{xy}.$ (41)
Here we have used Eq. (17) and (18), and further that $^{\bf w}=-w_{z}^{D}\hat{\bf z}$. Similarly, for the valleys $\bf K$$^{\prime}_{\pm}$ we find that the second term gives a contribution given by
$\partial_{x}{\varepsilon}^{{\bf K}^{\prime}_{\pm}}_{\bf q}\frac{{\bf w}\cdot{\bf n}^{{\bf K}^{\prime}_{\pm}}_{\bf q}\times\partial_{y}{\bf n}^{{\bf K}^{\prime}_{\pm}}_{\bf q}}{2\left|{\bf n}^{{\bf K}^{\prime}_{\pm}}_{\bf q}\right|^{3}}\rightarrow-w^{D}_{z}\frac{v_{0}}{v_{1}}\Omega^{{\bf K}^{\prime}_{\pm}}_{xy}$ (42)
Collecting all contributions and using (21) we then find the polarizability $\Lambda$ as
$\Lambda$ $=-\frac{e}{\pi\hbar}\left(w^{D}_{0}-\frac{v_{0}w^{D}_{z}}{v_{1}}\right)\text{sgn}(\Delta)$
$=-\frac{e}{\pi\hbar}\frac{\eta_{-}\delta(t^{2}_{1}-t^{2}_{0})}{t^{2}_{1}}\text{sgn}(\Delta).$ (43)
In the second line we have used Eqs. (38), (39), (26), (27), as well as the definitions of $v_{0}$ and $v_{1}$.
Appendix D Continuum Dirac quadrupole model of the Lieb lattice altermagnet
@@
We will proceed as follows. First, we determine the continuum Dirac model by expanding the lattice Hamiltonian in the absence of strain, and then perform a similar expansion of the strain terms, for which we only keep the lowest zeroth order contributions. Based on the linearized continuum model of the full Hamiltonian, we then show in Appendix 69 how the piezomagnetic polarizability is computed.
@@
These functions capture anisotropic next-nearest neighbor (and intra-sublattice) hoppings allowed in the presence of strain. In the continuum model expansion we only require the lowest order terms, which are constant $\bf q$}-independent terms obtained from evaluating $\chi_{\bf k}$ and $w_{z,{\bf k}}$ at $\bf K$$_{\pm}$ and $\bf K$$^{\prime}_{\pm}$. We find
@@
Appendix E Calculation of piezomagnetic polarizability for the Lieb lattice altermagnet
@@
Full expression of piezomagnetic polarizability. As stated in the main text, the piezomagnetic polarizability of the Lieb lattice model is a sum over the contributions from each spin sector, with each of these contributions being described by (4). The explicit expression for $\Lambda$ is then
$\Lambda=-\frac{e}{\hbar}\sum_{\sigma}\int\frac{d^{2}{\bf k}}{(2\pi)^{2}}\bigg\{\chi_{\bf k}\frac{{\bf n}^{\sigma}_{\bf k}\cdot\partial_{x}{\bf n}^{\sigma}_{\bf k}\times\partial_{y}{\bf n}^{\sigma}_{\bf k}}{2|{\bf n}^{\sigma}_{\bf k}|^{3}}\\ +\bigg[\partial_{x}\varepsilon_{\bf k}\frac{{\bf w}_{\bf k}\cdot{\bf n}^{\sigma}_{\bf k}\times\partial_{y}{\bf n}^{\sigma}_{\bf k}}{2|{\bf n}^{\sigma}_{\bf k}|^{3}}-(x\leftrightarrow y)\bigg]\bigg\},$ (69)
with ${\bf n}^{\sigma}_{\bf k}$ and $\varepsilon_{\bf k}$ defined in Eq. (5).
@@
For the valley $\bf K$$_{+}$, we can replace $\chi_{\bf k}$ by $w^{D}_{0}$ (see Appendix D) and we can take the vector ${\bf n}={\bf n}_{\bf q}$ as ${\bf n}=(-v_{1}q_{x},-\zeta m,v_{2}q_{y})$. The vector $\bf w$} is simply given by ${\bf w}=(0,0,w^{D}_{z})$. With these ingredients we evaluate both the geometric and the interband terms in (69) (i.e., the first and second terms in the integrand). From simple inspection we see that the geometric term is proportional to the Berry curvature of the occupied valence band, which is multiplied by the energy shift $w^{D}_{0}$. In the special case of the linearized Dirac model, the interband term can also be rewritten in terms of the Berry curvature (by multiplying and dividing by $v_{2}$), while $\partial_{y}\varepsilon_{{\bf K}_{+}+{\bf q}}$ becomes $\partial_{y}(-v_{0}q_{y})=-v_{0}$. Putting everything together, one obtains
@@
Appendix F Topological transition in the Lieb lattice altermagnet